OpenStudy (anonymous):

Check my answers please? And correct the ones that are wrong? 1. What is the sum? 5/3 + (-1/5) A. -22/5 B. -22/15 C. 22/15** D. 28/15 2. Which property is illustrated by the statement? (3+ 2.8)+5=3+(2.8+5) A. Commutative property of addition** B. Commutative property of multiplication C. Inverse property of multiplication D. Associative property of addition 3. To which subset of real numbers does the number 1/4 belong? A. Rational numbers** B. Irrational numbers C. Whole numbers, integers, rational numbers D. Whole numbers, natural numbers, integers I'm unsure of the last two

OpenStudy (blacksteel):

Your first answer is correct. For the second question: The commutative property of addition states that a + b = b + a The commutative property of multiplication states that ab = ba The inverse property of multiplication states that a * 1/a = 1 The associative property of addition states that (a + b) + c = a + (b + c) This particular statement demonstrates the associative property of addition. For the third querstion: Rational numbers are numbers that can be expressed as the ratio of two integers; ie., any number that can be written as a/b. Irrational numbers are numbers that cannot be expressed as the ratio of two integers. Whole numbers are integers greater than or equal to 0. Natural numbers are integers greater than 0. 1/4 is a rational number.

OpenStudy (anonymous):

Ok thank you :)